# Analyzing Fine-tuning Representation Shift for Multimodal LLMs Steering

### v1

@@ 4.2 Recovering concepts in fine-tuned model via shift vectors
To study if the fine-tuned concepts $\bm{U}^{b}$ can be recovered from the original ones $\bm{U}^{a}$, we first establish a matching ($m:i\rightarrow j$) between the set of original $\{\bm{u}^{a}_{i}\}_{i=1}^{K}$ and fine-tuned concepts $\{\bm{u}^{b}_{j}\}_{j=1}^{K}$. For systematic evaluation of recovery of all fine-tuned concepts, we constrain $m$ to be bijective (see App. A for more details). Finally, we evaluate how well a shifted concept $\bm{u}^{s}_{k}$ is similar to its match $\bm{u}^{b}_{m(k)}$ using the overlap metrics. Fig. 7 shows the results of recovering the fine-tuned concepts for models fine-tuned on different subsets of the VG dataset (place, color, sentiment). We compare the text grounding overlap between the shifted $u^{s}_{k}$ and fine-tuned concepts $u^{b}_{m(k)}$. We use the overlap between the original concepts $u^{a}_{k}$ with the fine-tuned ones as a baseline. As we can see, most shifted concepts exhibit higher overlap than the original ones. Many of them have very high overlap close to 100% (full recovery).
@@
Figure 7: Recovering fine-tuned model concepts. For each matched fine-tuned concept we report its text grounding overlap with the original concept and the shifted concept. Model fine-tuned on: places (top), colors (middle) and sentiments (bottom).
Shift magnitude ($\alpha$) and concepts recovery.
We also study the amount of recovery for shifted concepts, obtained with different shift magnitudes $\alpha$. We report the average recovery over $K=20$ concepts for each fine-tuning task for different $\alpha$ values in Fig. 8. Note that $\alpha=0$ corresponds to original concepts. $\alpha=1$ generally corresponds to the most optimal value of shift magnitude (color, sentiment fine-tuning) or very close to the optimal value (place fine-tuning). This indicates that simply adding the mean shift vector to the original concept without scaling, generally offers the best possibility to recover the fine-tuned concepts.
Figure 9: Correlation between shift vector magnitudes and concept recovery. Each point corresponds to an original concept from models fine-tuned on 3 VG subsets. The higher the shift vector magnitude the better the concept recovery.
Which concepts are recovered?
We examine the potential correlation between the computed concept shift vectors and the recovery of concepts in the fine-tuned model. We hypothesize that if samples related to an original concept consistently and significantly shift toward the corresponding fine-tuned model concepts, the shift vector, computed based on these samples, should be more effective at recovering this concept. Assuming that individual shifts $\delta_{m}^{a\rightarrow b}$ associated with each original concept tend to have similar magnitudes, their aggregation $\Delta_{k}^{a\to b}$ should have a higher magnitude when they are aligned. We validate that this assumption is reasonable for most concepts in our analysis (App. A).
For discovered concepts across all three fine-tuning tasks, we measure the correlation between the magnitude of the concept shift vectors $||\Delta_{k}^{a\to b}||$ and concepts recovery, here, measured as the cosine similarity between the shifted and matched fine-tuned concept. For a given shifted concept $\bm{u}^{s}_{k}$, this recovery (CR) is calculated as:
@@
This is illustrated in Fig. 9, where we observe a noticeable correlation between the two quantities, which further strengthens our hypothesis.
Figure 10: Steering MLLMs answers. Top: overview of the steering framework. We add a steering vector to the residual stream of MLLMs to change its response. Bottom: Each example corresponds to different steering vector that change a specific original answer (e.g., white) to a target one (e.g., black).
5 Fine-grained multimodal LLM steering
Model steering (see Fig. 10) refers to guiding the model outputs towards desired outcomes by modifying the features without altering the model weights. Here, we investigate this technique, focusing on large multimodal models for visual question-answering and image captioning tasks. We provide additional qualitative results and ablation study in App. B.
Motivation.
In the previous sections, we demonstrate the feasibility of recovering target concepts in fine-tuned models by shifting the original model features along shift vectors. In addition, we find that the features related to different concepts are almost linearly separable (Fig. 11), and this becomes more prominent in deeper layers. This validates the linear representation hypothesis for MLLMs, previously studied for LLMs [38, 32]. Finally, model steering might be an alternative method to avoid costly fine-tuning.
@@ 5 Fine-grained multimodal LLM steering
5.1 Multimodal LLMs steering framework.
Setup.
For a text-generative multimodal model $f:T\times I\rightarrow T$, we aim to modify the model’s output $\hat{y}$ to a desired output $\tilde{\bm{y}}$ by applying a shift or steering vector $\bm{v}$ to the residual stream features $\bm{Z}$, without changing the model parameters. The model is evaluated within a visual question-answering (VQA) framework: where each query consists of a question about an input image, and the model generates an answer. This task encompasses a wide range of applications for MLLMs. To measure the effectiveness of our approach in directing the model towards specific answers or answer types, we report the number of generated answers that align with the target output or answer type. Additionally, we aim for targeted steering, ensuring that only specific answer types are influenced. For example, when altering answers from “yes” to “no” within the ”yes/no” category, responses to other question types should remain unaffected. This specificity is assessed by monitoring accuracy across answer types and tracking the number of answers from each answer type.
Implementation details.
Experiments are conducted on the widely-used LLaVA model [30], comprising a CLIP image encoder, a two-layer MLP connector, and a 7B Vicuna-1.5 LLM. In the main paper, we focus on VQAv2 dataset [18], a visual question-answering corpus with image-question-answer triplets and annotated answer types (”yes/no”, ”number”, and ”other”). We provide also experiments on COCO captioning [29], which we detail more in App. B. Steering vectors are derived from a subset of the training set, with model performance evaluated on the validation set. Additional experiments and ablation study can be found in App. B. We find that the steering becomes more effective in deeper layers (see Fig. 11 and App. B), thus we apply it on the last layer for VQAv2.
@@
Table 1: Steering MLLMs answers type. Number of target answers type increases significantly after steering the model with the corresponding steering vector.
5.2 Coarse-grained model steering
In coarse-grained or global steering, the objective is to adjust the model outputs $\hat{y}$ to generally align with a set of target samples (e.g., changing answers type). Given input-output samples, we first extract the answer representations $\bm{B}={\bm{b}_{1},...,\bm{b}_{N}}$ at a specified layer $l$ (we drop the layer index for simplicity) from the target sample set. Similarly, we obtain representations for a set of original samples $\bm{A}={\bm{a}_{1},...,\bm{a}_{M}}$, for example, randomly drawn from the training set. We then compute the coarse steering vector $\bm{s}_{c}$ as follows:
$\bm{s}^{c}=\frac{\sum_{i}^{N}\bm{b}_{i}}{N}-\frac{\sum_{i}^{M}\bm{a}_{i}}{M},$ (7)
$s_{c}$ is applied to all the samples in the validation set. For instance, the activations $z_{i}$ of a sample $x_{i}$ (at the same layer $l$), are changed as follows:
$\tilde{\bm{z}}_{i}=\bm{z}_{i}+\alpha\bm{s}_{c},\quad\bm{z}_{i}=f_{l}(x_{i}),$ (8)
where $\alpha$ controls the steering strength and it is set to 1 (we study $\alpha$ in App. B). $\tilde{\bm{z}}$, replaces $\bm{z}$ and becomes the input to the next layer $l+1$.
Experiments.
We steer the model answers towards a particular type among; yes/no, numbers or other (e.g. colors, objects). For each target answer type, we compute a steering vector. Table 1 shows that the number of answer types increases significantly when applying the corresponding steering vector, this validates the efficacy of globally steering the model.
Figure 12: Discovering meaningful steering directions. Each line corresponds to a fine-grained steering direction to steer the model answer to: ”No” (yes/no), ”4” (number) and ”Red” (other). Some steering directions are targeted (e.g., ”No”) as there is slight change in both the accuracy on other types (e.g., number, other) and the number of answers type. We show the relative scores compared to a baseline with no steering.
5.3 Discovering meaningful and fine-grained steering directions.
Unlike global steering, fine-grained steering focuses on adjusting outputs at the concept level. Specifically, we seek editing directions that adjust only certain concepts. To do this, we decompose the hidden states of a set of samples into a set of concepts $\bm{U}$. We then compute a series of fine-grained steering vectors $\bm{s}^{f}={\bm{s}_{11}^{f},...,\bm{s}_{NN}^{f}}$, where each $\bm{s}_{ij}^{f}$ represents the steering vector from concept $\bm{u}_{i}$ to $\bm{u}_{j}$:
$\bm{s}_{ij}^{f}=\bm{u}_{j}-\bm{u}_{i}.$ (9)
However, not all computed vectors are necessarily meaningful steering vectors. We identify these vectors based on the strongest impact on guiding the model towards generating specific answers or concepts (e.g. producing significantly more target answers). This is more detailed in App. B.
Experiments.
Fig. 12 illustrates some steering vectors. We try to find the steering vectors between concepts decomposed from 3 sample sets corresponding to: ”yes/no”, ”number” and ”other” answers type. We notice that some vectors correspond to answering ”No”, ”4” and ”Red”. This validates the possibility of finding fine-grained steering vectors that can steer the model towards a specific answer.
@@ 5.2 Coarse-grained model steering
Table 2: Steering MLLMs answers. Steering answers from ”Yes” (Yes/No), ”1” (Number), ”White” (Other) to ”No”, ”3”, ”Black” respectively. The number of original/target answer counts decreases/increases significantly, while the accuracy on other answer types changes slightly, and the number of answer type counts remains almost constant.
5.4 Steering towards a specific target answer
Motivated by the previous section, the objective here is to steer the model towards a specific answer specified by the user. For each pair of original/target answers (e.g., yes/no), we collect few hundred samples and compute the steering vector (as in Section 5.2). Then we apply the vectors on all samples in the validation set. In Table 2, we report the evaluation metrics when targeting the last layer. We can notice that the steering is effective as the number of target answers increase, while the accuracy on other answer types only slightly changes.
@@ 5.4 Steering towards a specific target answer
5.5 Steering image captions
In our earlier experiments, we applied steering on relatively brief answers from the VQAv2 dataset. Here, we extend this approach to longer, descriptive outputs using the COCO image captioning dataset [29]. Given that multiple captions can effectively describe an image by emphasizing various aspects such as the main object, surroundings, actions, or events, our focus is modifying captions to align with a specific target style (see Fig. 13). We presents the results in Table 3, demonstrating that captions can be steered towards a target style so that they focus more on places, colors, or sentiments. These findings validate the feasibility of steering MLLMs on tasks with longer responses.
@@ 5.5 Steering image captions
Table 3: Steering MLLMs captions style. Each line corresponds to different steering vector. Steering towards a target style increases the number of captions with that style.
@@ 6 Discussion
Ideally, steering should be targeted; shifting from a “yes” to a “no” should affect only questions initially answered with “yes”. While this is often achieved, it is not perfect. We also notice a tradeoff between the steering strength and the quality of generated responses. This approach focuses on changing the model answers or style, however, it can be extended to other applications, such as addressing biases or mitigating safety concerns. It is also interesting to extend our study to larger and more recent MLLMs, eventually with different architectures.
Conclusion.
Gaining insights into the mechanisms of recent foundation models is crucial for advancing AI research. In this study, we delve into how MLLMs’ internal representations evolve during fine-tuning, demonstrating that concepts learned post-fine-tuning can often be recovered from the original model by navigating the feature space. This approach inspired our final investigation, where we steer model behavior by modifying features directly, without additional training. We show that this could be used to change model answers, or have richer captions focusing on different aspects of the image. We hope this work motivates further research into understanding these models and developing more efficient methods for adapting or refining them, moving beyond traditional training practices.
@@ Conclusion.
The code for this project is publicly available at https://github.com/mshukor/xl-vlms.
@@ Appendix A Fine-tuning and evolution of concept representations
This section provides additional details and analyses on the evolution of concepts due to fine-tuning and their recovery using shift vectors. Section A.1 introduces additional notations. Section A.2 describes our experiments’ models, fine-tuning setup, and datasets. Section A.3 analyzes the change of concepts during training. In Section A.4.2, we present ablation studies related to concepts recovery. Section A.5 discusses the correlation between the concepts recovery and the consistency of their shifts. Finally, Section A.6 covers experiments conducted with the LLaVA-1.5 [30].
@@ A.1 Notations
In the main paper, we focus on the representation in the residual stream. Similar to [12], this can be expressed as follows:
@@ A.2 Implementation details
Our analysis spans MLLMs following the architecture detailed in the paper. We distinguish 2 setups; single-task tuning that covers most of the analysis part in the main paper, and multi-task tuning with additional results in the appendix. For multi-task setup, we use LLaVA [30], that consists of a CLIP image encoder, a two-layer MLP connector, and a 7B Vicuna-1.5 LLM. For single-task setup, we follow the setup in [37, 54, 45].
We fine-tune the LLM with Low-Rank Adaptation (LoRA) [19], which modifies the weight matrices of the model with a low-rank update. We use AdamW optimizer with a weight decay of $0.01$ and choose the learning rate and LoRA rank that works best for each fine-tuning dataset. For LLaVA, we follow the hyperparameters recommended by the authors, including the rank $r=128$ and learning rate $2\mathrm{e}{-4}$.
We fine-tune the models using three distinct subsets of Visual Genome (VG) dataset [24]: color, sentiment, and place. These subsets respectively correspond to about $21k$ samples describing colors, $5k$ samples containing sentiments and $27k$ samples that describe the locations or environments. All subsets were curated based on keyword occurrences provided in Fig. 15. We also use COCO captioning dataset [29] for hidden states extraction, throughout the quantitative experiments. Different than VG, COCO contains captions describing the image general, often focusing on the central object.

### v2

@@ 4.1 Fine-tuning experiments
Evaluating fine-tuned concept recovery. To study if the fine-tuned concepts $\bm{U}^{b}$ can be recovered from the original ones $\bm{U}^{a}$, we first establish a matching ($m:i\rightarrow j$) between the set of original $\{\bm{u}^{a}_{i}\}_{i=1}^{K}$ and fine-tuned concepts $\{\bm{u}^{b}_{j}\}_{j=1}^{K}$. For systematic evaluation of recovery of all fine-tuned concepts, we constrain $m$ to be bijective using an optimal transport algorithm detailed in Section A.1. Finally, we evaluate how well a shifted concept $\bm{u}^{s}_{k}$ (Equ. (4)) is similar to its match $\bm{u}^{b}_{m(k)}$ using the aforementioned T-Overlap metric. Fig. 6 shows the results of recovering the fine-tuned concepts for models fine-tuned on different subsets of the VG dataset (place, color, sentiment). We report the T-Overlap between the shifted $u^{s}_{k}$ and fine-tuned concepts $u^{b}_{m(k)}$ for various different tokens of interest. For each target token and finetuning we extract $K=20$ concepts and report the mean and standard deviation over them We use the overlap between the original concepts $u^{a}_{k}$ and the fine-tuned ones as a baseline. We observe that most shifted concepts show higher overlap than the original ones : this demonstrates that fine-tuned concepts can be efficiently recovered from the original model with concept shift vectors.
@@
Figure 6: Recovering fine-tuned concepts. Across different fine-tunings (places (top), colors (middle) and sentiments (bottom)), we compute the average text grounding overlap of original and shifted concepts with the matched fine-tuned ones. Shifting the original concepts result in partially recovering the fine-tuned ones.
Figure 7: Correlation between shift consistency and concept recovery (Color finetuning). The more consistent and aligned the individual representation shifts associated with a concept, the better the recovery of the fine-tuned concept.
Which concepts are recovered better? We hypothesize that if the representation shift for individual samples associated with a concept $\bm{u}_{k}^{a}$, $\{\delta^{a\to b}_{m}\;|m\in\bm{A}_{k}\}$, is consistently aligned with the concept shift vector, the resulting $\bm{\Delta}_{k}^{a\to b}(\bm{u}^{a}_{k})$, should be more effective at recovering the fine-tuned concept. We quantify the consistency through mean cosine similarity of $\{\delta^{a\to b}_{m}\}_{m\in\bm{A}_{k}}$ with the concept shift vector $\bm{\Delta}_{k}^{a\to b}(\bm{u}^{a}_{k})$. In other words, this quantifies the alignment of individual shifts and their mean, $\bm{\Delta}_{k}^{a\to b}(\bm{u}^{a}_{k})$:
We measure the recovery of concept $k$, $\text{CR}_{k}$, as the improvement in similarity between the matched fine-tuned $\bm{u}^{b}_{m(k)}$ and shifted concept $\bm{u}^{s}_{k}$, relative to the original one $\bm{u}^{a}_{k}$:
@@
We plot the consistency and recovery for concepts extracted across four tokens of interest for color finetuning in Fig. 7. Plots for other finetunings are in Section A.5. Crucially, we observe a positive and statistically significant correlation between the two quantities across all the finetuning tasks. This supports our hypothesis that a better concept shift recovery is related to consistency of individual shifts of original concept. More ablation studies about concept recovery are available in Section A.4, analyzing the influence of steering strength $\alpha$, number of concepts $K$, and extraction layer $l$.
@@
In summary, we demonstrated the feasibility of recovering target fine-tuned concepts by applying simple per-concept shift of the original model features. This supposes that the features related to different concepts are almost linearly separable, which empirically seems to hold at least for the last MLLM layers as pictured on Fig. 8, and as previously studied for LLMs in [48, 42]. This motivates the following investigations on using a similar methodology for simple, computationally inexpensive, yet efficient MLLM steering.
4.2 Multimodal model steering
We perform steering by applying a steering vector $\bm{v}$ to the residual stream features $\bm{Z}$ of the MLLM, without changing its parameters. We first evaluate the MLLM steering capabilities in a visual question-answering (VQA) setup. Then, we show the applicability of our MLLM steering to control captioning styles. Lastly, we present two steering applications: gender debiasing, aiming to mitigate biases in model outputs, and safety alignment, i.e. ensuring that the model refuses to generate harmful information. We discuss technical details for each of these applications in App. C and App. D.
Setup. For VQA tasks, each query consists of a question about an input image, and the model generates an answer. To measure the effectiveness of our approach in directing the model towards specific answers or answer types, we report the number of generated answers that align with the target output or answer type. Additionally, we aim for targeted steering, ensuring that only specific answer types are influenced. For example, when altering answers from “yes” to “no” within the ”yes/no” category, responses to other question types should remain unaffected. This specificity is assessed by tracking accuracy across answer types and number of answers from each type.
Implementation details. Experiments in the main paper are primarily conducted on LLaVA [39] for conciseness. However, we show in Section B.2 that our method is generally applicable to other popular MLLMs. We experiment on VQAv2 [24], a visual question-answering corpus with image-question-answer triplets and annotated answer types (”yes/no”, ”number”, and ”other”). Steering vectors are derived from a subset of the train set, with model performance evaluated on the val set. As steering becomes more effective in deeper layers (see Fig. 8 and App. B), we apply it on the last layer. Additional experiments can be found in App. B.
@@ 4.2 Multimodal model steering
Table 1: Steering MLLMs answers type. Number of target answers type increases significantly after model steering.
Figure 9: Discovering meaningful steering directions. Each line corresponds to a fine-grained steering direction to steer the model answer to: ”No” (yes/no), ”4” (number) and ”Red” (other). Some steering directions are targeted (e.g., ”No”) as there is slight change in both the accuracy and number of answers types on other types (e.g., number, other). We show the relative scores compared to a baseline with no steering.
Coarse and Fine-Grained Model Steering for VQA. We explore both coarse and fine-grained steering in VQA tasks. Specifically, coarse steering aims to alter the distribution of answers, while fine-grained steering targets specific responses. For coarse-grained steering, we direct the model’s answers toward a particular category: yes/no, numbers, or other (e.g., colors, objects). We compute a steering vector for each target answer type. As shown in Table 1, applying steering significantly increases the proportion of the targeted answer type. For fine-grained steering, we first assess the feasibility of identifying such steering vectors. Fig. 9 illustrates examples of these vectors. We derive them from three sets of sample answers corresponding to ”yes/no,” ”number,” and ”other” categories. Interestingly, some vectors distinctly align with specific answers such as ”No,” ”4,” and ”Red.” This confirms the potential to identify fine-grained steering vectors capable of guiding the model toward a precise response. Building on these insights, we seek to steer the model toward a user-specified answer. For each original/target answer pair (e.g., yes/no), we collect some samples and compute the corresponding (coarse) steering vector. We then apply these vectors to all validation set samples. In Table 2, we report evaluation metrics when steering at the last layer. The results show that steering effectively increases the occurrence of target answers, while accuracy on other answer types remains largely stable.
@@
Table 2: Steering MLLMs answers. Steering answers from ”Yes” (Yes/No), ”1” (Number), ”White” (Other) to ”No”, ”3”, ”Black” respectively. The number of original/target answer counts decreases/increases significantly, while the accuracy on other answer types changes only slightly, and the number of answer type counts remains almost constant.
@@
Steering image caption styles. We previously applied steering on relatively brief answers from the VQAv2 dataset. Here, we extend this approach to longer, descriptive outputs using the COCO captioning dataset [37]. Given that multiple captions can effectively describe an image by emphasizing various aspects such as the main object, surroundings, actions, or events, we aim at modifying captions to align with a specific target style. Here, we learn a coarse steering vector between samples with predicted captions in the target style and random samples. Qualitative examples of this steering for LLaVA are in Fig. 10. Results in Table 3 demonstrate that captions can be effectively steered towards a target style even when considering tasks with longer responses. We provide more captioning experiments in Section B.4.
@@
Table 3: Steering MLLMs captions style. Each line corresponds to a different steering vector. Steering towards a target style increases the number of captions with that style.
Gender debiasing captions.
We perform gender debiasing on COCO test set, aiming at mitigating biases for any gendered nouns when captioning. We experiment with both coarse and fine-grained steering. The coarse steering vector is computed between sets of samples with a gendered/neutral noun in the caption. For fine-grained steering, we steer each concept to its closest (as defined by its cosing similarity) counterpart among the neutral concepts, as explained in Section 3.3. The results are reported in Table 4. Interestingly, both strategies are capable of converting many gendered captions to neutral ones, with fine-grained steering being significantly more effective than coarse steering.
Model Total Method Gendered $\rightarrow$ Neutral
LLaVA-1.5 794 coarse 232
fine-grained 632
Idefics2 815 coarse 237
fine-grained 315
Qwen2-VL-Instruct 926 coarse 134
fine-grained 300
Before Steering After Steering
A young boy with curly hair is playing a video game. A child with curly hair is playing a video game.
A man riding a dirt bike on a beach. A person riding a dirt bike on a beach.
Table 4: Gender debiasing results: number of occurrences of gendered terms converted to neutral terms across different models and methods, after steering with $\alpha=1$. Below, qualitative samples illustrate changes in descriptions before and after applying steering.
Safety alignment. Pure text LLMs often exhibit stronger safety alignment compared to MLLMs [11]. Empirical evidence for this can be found in App. D. Using this insight, we construct two sets of samples, categorized as safe and unsafe, by evaluating the model’s response to identical malicious content presented in different modalities: one conveyed through text and the other through text + image. We use these to compute a safety guard steering vector. This steering vector gives the model a higher level of safety, without affecting its usefulness for safe tasks (Table 5). We evaluate the safety of the model by ASR (attack success rate) metric, described in detail in App. D.
Model Before steering After steering
Qwen2-VL-Instruct 45/100 5/100
Table 5: Enhancing MLLM Safety Through Steering. We evaluate safety using the ASR metric, quantifying the proportion of responses that do notrefuse to provide harmful instructions. Our assessment is conducted on a portion of MM-SafetyBench [40] dataset. A lower ASR is desired when the prompt requires harmful instructions.
@@ 5 Discussion
The effectiveness of our method relies on the fact that the concepts are represented by linear directions in latent space. However, recent work [14] has found that not all features are captured as such : thus, when analyzing recovery of fine-tuned concepts, it can be interesting to explore more sophisticated similarity measures and matching algorithms.
Conclusion. In this work, we introduced a novel concept-based analysis framework for monitoring and controlling MLLM behavior, offering new insights into how latent representations evolve during fine-tuning and across datasets. To address the former, we proposed concept shift vectors, an efficient method for recovering and interpreting concepts in fine-tuned models relative to their original counterparts. This approach naturally led us to explore the latter case, where we demonstrated the ability to steer model behavior by modifying features – without requiring additional training. Our results show that this technique effectively modifies MLLM answers, enabling applications such as gender debiasing, safety control, and enhanced caption generation that highlight different aspects of an image. By releasing our code, we hope our framework will benefit the community and encourage research towards better understanding of MLLMs, as well as their broader applications in domains such as physical and digital agents [60, 50].
Acknowledgements
This work has been partially supported by ANR grant VISA DEEP (ANR-20-CHIA-0022), HPC resources of IDRIS under the file A0160614966 allocated by GENCI, and Cluster PostGenAI@Paris (ANR-23-IACL-0007, FRANCE 2030).
@@ Acknowledgements
  • App. C details our experiments for gender debiasing.
  • App. D includes additional details and results to steer for safety.
@@ Appendix A Fine-tuning and evolution of concept representations
This section provides additional details and analyses on the evolution of concepts due to fine-tuning and their recovery using shift vectors. Section A.1 introduces additional notations. Section A.2 describes our experiments’ models, fine-tuning setup, and datasets. Section A.3 analyzes the change of concepts during training. In Section A.4, we present ablation studies related to concepts recovery. Section A.5 discusses the correlation between the concepts recovery and the consistency of their shifts.
@@ A.1 Notations
In this paper, we particularly focus on the representations in the residual stream [13]. This can be expressed as follows:
@@ A.2 Implementation details
Our analysis spans MLLMs following the architecture detailed in the paper. We distinguish 2 setups; multi-task tuning (main paper), and single-task tuning with additional results in the appendix. For multi-task setup, we use LLaVA [39], that consists of a CLIP image encoder, a two-layer MLP connector, and a 7B Vicuna-1.5 LLM. For single-task setup, we follow the setup in [47, 67, 56].
We fine-tune the LLM with Low-Rank Adaptation (LoRA) [25], which modifies the weight matrices of the model with a low-rank update. We use AdamW optimizer with a weight decay of 0.010.010.01 and choose the learning rate and LoRA rank that works best for each fine-tuning dataset. For LLaVA, we follow the hyperparameters recommended by the authors, including the rank $r=128$ and learning rate $2\mathrm{e}{-4}$.
[…]